# PolygloToxicityPrompts[Uncaptioned image] : Multilingual Evaluation of Neural Toxic Degeneration in Large Language Models

### v2

@@ 1 Introduction
We introduce PolygloToxicityPrompts (PTP), [^1] We provide our dataset, code, and a leaderboard: https://huggingface.co/spaces/ToxicityPrompts/PTP the first large-scale multilingual benchmark for evaluating neural toxic degeneration, defined as the propensity of LLMs to generate toxic text given a prompt (Gehman et al., 2020). We create PTP by scraping over 100M documents from web-text corpora to collect naturally occurring toxic prompts. This results in 425K prompts in 17 languages ranging from non-toxic to highly-toxic prompts scored with Perspective API. [^2] https://perspectiveapi.com/
@@ 3.1 Operationalizing and Evaluating Toxicity
We define toxicity as “a rude, disrespectful, or unreasonable comment that is likely to make people leave a discussion” (Wulczyn et al., 2017; Borkan et al., 2019). We use Perspective API,2 an industry-standard toxicity detection tool because it supports our 17 languages. Specifically, we use the Toxicity score from Perspective API, computed using the UTC (Unified Toxic Content Classification) framework (Lees et al., 2022), composed of a Charformer-based transformer (Tay et al., 2022). UTC is a Seq2Seq architecture pretrained with the mC4 corpus (Xue et al., 2021) and Perspective Pretraining Corpus (PPC). Additionally, Perspective API utilizes a single-language CNN (Lecun et al., 1998) distilled from multilingual BERT models (Devlin et al., 2019) for German and Portuguese.
@@ 3.2 Dataset Creation
The final dataset includes 25K naturally occurring prompts for each language, for a total of 425K prompts across 17 languages. Figures 10(a) and 10(b) show the prompt toxicity and length distributions of our collected prompts for all languages. We create our prompts using documents instead of sentences (Gehman et al., 2020). Thus, our prompts are much longer than RealToxicityPrompts, with an average length of approximately 400 GPT-4 tokens.
@@
While the extraction of toxic content from web-text may appear straightforward, we encountered several challenges associated with the scarcity of multilingual toxicity. The mC4 corpus (Xue et al., 2021) filters toxicity by removing pages containing bad words.4 As a result, we observe less than $0.01\%$ toxicity rate out of 5M samples for ar, cs, fr, ko, id, it, nl, pl, and sv. However, consistent with previous findings (Zhou et al., 2021; Dodge et al., 2021), we note that filtered datasets still exhibit toxicity, and observe higher toxicity rates for other languages.
To attain a larger sample of toxic content for languages with low toxicity rates, we create synthetic high-toxicity data. Specifically, we translate toxic samples from the mC4 and The Pile corpora into target languages using the NLLB-3B model (Team et al., 2022). We use this process to create $\approx$ 70K translated prompts across 9 languages, which amounts to only $16.8\%$ of our dataset. Contrary to prior works, we observe a Pearson correlation of 0.725 ($p\leq 0.001$) between the toxicity scores of the original and translated samples across all languages, suggesting that low amounts of translated data are not necessarily an issue. [^5] We discuss limitations with translating data in the Ethics Statement.
@@
We also create $\text{PTP}_{\text{Small}}$, a stratified sample of 5K prompts per language from PolygloToxicityPrompts to benchmark models in limited computational resources.
@@ 4.1 How does Prompt Language impact Average Toxicity?
Figure 3: Language-wise AT trends for multilingual models. Takeaway: High toxicity scores for all languages indicate the need for multilingual toxicity mitigation methods.
@@ 4.2 How does Model Size impact Average Toxicity?
Figure 6: AT for differet model categories. Takeaway: base $>$ instruct $\approx$ preference.
@@ A.2 Dataset Statistics
Figures 10(a) and 10(b) show the Toxicity score and length distribution for prompts in PolygloToxicityPrompts. We calculate prompt length in terms of GPT-4 tokens using tiktoken [^16] https://github.com/openai/tiktoken.
(a) Prompt toxicity score distribution
(b) Prompt length distribution
Figure 10: Distributions of prompt toxicity and prompt length for PTP

### v3

@@ Abstract
[Uncaptioned image] Code kpriyanshu256/polyglo-toxicity-prompts
[Uncaptioned image] Dataset ToxicityPrompts/PolygloToxicityPrompts
[Uncaptioned image] Leaderboard ToxicityPrompts/PTP
@@ 1 Introduction
We introduce PolygloToxicityPrompts (PTP), the first large-scale multilingual benchmark for evaluating neural toxic degeneration, defined as the propensity of LLMs to generate toxic text given a prompt (Gehman et al., 2020). We create PTP by scraping over 100M documents from web-text corpora to collect naturally occurring toxic prompts. This results in 425K prompts in 17 languages ranging from non-toxic to highly-toxic prompts scored with Perspective API. [^1] https://perspectiveapi.com/
@@ 3.1 Operationalizing and Evaluating Toxicity
We define toxicity as “a rude, disrespectful, or unreasonable comment that is likely to make people leave a discussion” (Wulczyn et al., 2017; Borkan et al., 2019). We use Perspective API,1 an industry-standard toxicity detection tool because it supports our 17 languages. Specifically, we use the Toxicity score from Perspective API, computed using the UTC (Unified Toxic Content Classification) framework (Lees et al., 2022), composed of a Charformer-based transformer (Tay et al., 2022). UTC is a Seq2Seq architecture pretrained with the mC4 corpus (Xue et al., 2021) and Perspective Pretraining Corpus (PPC). Additionally, Perspective API utilizes a single-language CNN (Lecun et al., 1998) distilled from multilingual BERT models (Devlin et al., 2019) for German and Portuguese.
@@ 3.2 Dataset Creation
The final dataset includes 25K naturally occurring prompts for each language, for a total of 425K prompts across 17 languages. Figures 10(a) and 11(a) show the prompt toxicity and length distributions of our prompts for all languages. We create our prompts using documents instead of sentences (Gehman et al., 2020). Thus, our prompts are much longer than RealToxicityPrompts, with an average length of approximately 400 GPT-4 tokens (cl100k_base tokenizer).
@@
While the extraction of toxic content from web-text may appear straightforward, we encountered several challenges associated with the scarcity of multilingual toxicity. The mC4 corpus (Xue et al., 2021) filters toxicity by removing pages containing bad words.3 As a result, we observe less than $0.01\%$ toxicity rate out of 5M samples for ar, cs, fr, ko, id, it, nl, pl, and sv. However, consistent with previous findings (Zhou et al., 2021; Dodge et al., 2021), we note that filtered datasets still exhibit toxicity, and observe higher toxicity rates for other languages.
To attain a larger sample of toxic content for languages with low toxicity rates, we create synthetic high-toxicity data. Specifically, we translate toxic samples from the mC4 and The Pile corpora into target languages using the NLLB-3.3B model (Team et al., 2022). We use this process to create $\approx$ 70K translated prompts across 9 languages, which amounts to only $16.8\%$ of our dataset. Contrary to prior works, we observe a Pearson correlation of 0.725 ($p\leq 0.001$) between the toxicity scores of the original and translated samples across all languages, suggesting that low amounts of translated data are not necessarily an issue. [^4] We discuss limitations with translating data in the Ethics Statement.
@@
We also create $\text{PTP}_{\text{Small}}$, a stratified sample of 5K prompts per language from PolygloToxicityPrompts to benchmark models with limited computational resources.
@@ 4.1 How does Prompt Language impact Average Toxicity?
Figure 3: Language-wise AT trends for multilingual models. Takeaway: High toxicity scores (relative to the AT levels shown in Figure 1 and Table 1) for all languages indicate the need for multilingual toxicity mitigation methods.
@@ 4.2 How does Model Size impact Average Toxicity?
Figure 6: AT for different model categories. Takeaway: base $>$ instruct $\approx$ preference.
@@ A.2 Dataset Statistics
Figure 10 shows the distribution of scores for Toxicity attributes computed by Perspective API, namely, Toxicity score, Insult score, Threat score, Profanity score, Identity Attack score, and Severe Toxicity score for prompts in PolygloToxicityPrompts. We observe a relatively higher amount of toxicity related to the Insult and Profanity categories as compared to the other categories.
We calculate prompt length in terms of GPT-4 tokens (Figure 11(a)) using tiktoken [^15] https://github.com/openai/tiktoken.
(a) Toxicity score distribution
(b) Insult score distribution
(c) Threat score distribution
(d) Profanity score distribution
(e) Identity Attack score distribution
(f) Severe Toxicity score distribution
Figure 10: Distributions of scores across toxicity attributes computed by Perspective API for PTP.
(a) Prompt length distribution
(b) Llama Guard score distribution
Figure 11: Distributions of prompt length and Llama Guard score for PTP
For prompts in the English split of PTP, we compute the Llama Guard Scores (Figure 11(b)). The distribution is similar to the distribution of the toxicity scores (Figure 10(a)). We also tabulate the categories violated by the prompts as generated by the model in Table 3, where most of the unsafe prompts belong to the Sexual Content category.
Category Count
Safe 16551
Violence and Hate 351
Sexual Content 7823
Criminal Planning 60
Guns and Illegal Weapons 1
Regulated or Controlled Substances 32
Self-Harm 12
Table 3: Distribution of safety categories for PTP English split